HUBBLE SPACE TELESCOPE DAILY REPORT #2560 PERIOD COVERED: 0000Z (UTC) 02/14/00 - 0000Z (UTC) 02/15/00 Daily Status Report as of 046/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ0751+2 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. As documented in HSTAR 7505 and described in 2.1, due to the acquisition problems, this proposal may have been affected. Otherwise, the proposal completed with no further eported problems. 1.3 Completed WF/PC-2 8436 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to make archive observations parallel to the prime science observations as part of a POMS test proposal. This is a POMS test proposal designed to simulate scientific plans. As documented in HSTAR 7505 and described in 2.1, due to the acquisition problems, this proposal may have been affected. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.4 Completed Two Sets of STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. As documented in HSTAR 7505 and described in 2.1, due to the acquisition problems, the first iteration of this proposal may have been affected. Otherwise, the proposal completed nominally. 1.5 Completed WF/PC-2 8338 (The Evolution of the Host Galaxies of Powerful Radio Sources) The WF/PC-2 was used to make observations of the host galaxy of the powerful radio source known as MRC0529-21. A combined data set of radio galaxies and quasars will allow powerful tests of the relationship between radio galaxies and quasars and the evolution of the host galaxies of powerful radio sources over a wide range of look-back times. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.7 Completed STIS/CCD 8412 (STIS CCD Imaging Flats C8) The Space Telescope Imaging Spectrograph (CCD) was used to investigate the flat-field stability of the unit over a monthly period. A bias calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD 8411 (STIS CCD Spectroscopic Flats C-8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectroscopic mode. The observations were completed as scheduled, and no anomalies were reported. 1.9 Completed Two Sets of STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ1451-2 and HCLQ0941-1 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of WF/PC-2 8442 (WF/PC-2 Cycle 8 Standard Darks) The WF/PC-2 was used to make a series of Cycle 8 standard dark calibration observations. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Six Sets of FGS 8524 (FGS-2R Mini OFAD) FGS-2R was used to calibrate FGS2r's Optical Field Angle Distortion {OFAD} to a level of precision required to operate the FGS2r as a guiding instrument. HST will be pointed so as to cause selected stars in an astrometric cluster to be distributed throughout the FGS2r full FOV or pickle. The apparent relative positions of these stars will be measured by a sequence of Position Mode observations. In each of 5 orbits, each with a different pointing, about 15 stars will be observed. Their apparent relative positions will be compared to the astrometric catalog and the OFAD will be calibrated to remove such differences. The observations completed with no reported problems. 1.12 Completed Seven Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.13 Completed Six Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Two Sets of STIS/MA1/MA2 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.15 Completed Four Sets of FGS 8523 (FGS2R - FGS Alignment) FGS-2R was used to determine the alignment of the newly installed FGS-2R with FGS-1R and FGS-3. After the final adjustment of FGS-2R's AMA {proposal 8522}, selected stars from the astrometric M35 cluster will be observed by FGS2r in Position Mode, while FGS-1R and FGS-3 will lock onto astrometric guide star pairs. The pointing and orientation of HST will be held constant for the four visits of this test. Each visit will use different guide star pairs, but the same set of stars will be observed in FGS-2R. The observations completed with no reported problems. 1.16 Completed STIS/CCD/MA1 8107 (What is the Nature of the Cold Medium in Cooling Flow Clusters?) The Space Telescope Imaging Spectrograph (CCD/MA1) was used to observe X-ray spectra of cooling flow clusters of galaxies in an attempt to discover the presence of large columns of "cold'' absorbing material. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 14 Successful: 14 HSTAR 7505 reports that the acquisition at 045/004902Z acquired successfully, but at 045/011519Z an FGS acquisition command caused stop flags on FGS-3. FGS-1 then failed the search radius exceed limit. The proposals described in 1.2, 1.3 and 1.4 may have been affected. All this occurred as a result of reloading FGS K factor guiding defaults via the SMS. Scheduled Re-acquisitions: 2 Successful: 2 2.2 FHST Updates: Scheduled: 24 Successful: 24 2.3 Operations Notes: Servicing Mission Observatory Verification (SMOV) continues with the operations team still using CCS string "G". Using ROP DF-18A, the ESB was dumped at cleared at 045/1820Z. An SI C&DH error was reset at 045/2149Z per ROP NS-5. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW